OpenStudy (anonymous):

CAN SOMEONE PLEASE CHECK MY ANSWERS! I WILL GIVE U AS MANY MEDALS AS I CAN AND FAN U!!!!!! simplify the sum 1.(3x^4+2x^2+1)+(5x^4-9x-7) A.8x^4+2x^2-9x-6 B.-2x^4+2x^2+9x+8 C.8x^4-7x^2-9x-6 ** D. 8x^4+2x^2-9x-6 2. find the GCF of the term of the polynomial 42x^5+60x^4 A.42x^4** B.x^4 C.6x^6 D.6x^4 simplify the product using FOIL (4x-3)(3x+5) A.12x^2+11x-15 B.12x^2+11x+15 C.12x^2-29x-15 D.12x^2-29x+15** what is the simpler form of the expression (3h+2)(2h^2-7h+1) A.6h^3-17h^2+11h+3 B.6h^3-25h^2-17h+2** C.6h^3-17h^2-11h+2 D.6h^3-17h^2-11h+3

OpenStudy (texaschic101):

#1 is wrong #2 is wrong #3 is wrong #4 is wrong

OpenStudy (anonymous):

well can u help me?

OpenStudy (texaschic101):

1. 3x^4 + 2x^2 + 1 + 5x^4 - 9x - 7 combine like terms 3x^4 + 5x^4 = 8x^4 2x^2 -- no other like term -9x -- no other like term -7 + 1 = -6 put it all together and you get : 8x^4 + 2x^2 - 9x - 6

OpenStudy (texaschic101):

2. 42x^5 + 60x^4 = 6x^4 (7x + 10) 6x^4 is the greatest common denominator

OpenStudy (texaschic101):

3. (4x - 3)(3x + 5) ---- use FOIL first -- multiply first terms in each set ... 4x * 3x = 12x^2 outside -- multiply outside terms in each set ... 4x * 5 = 20x inside -- multiply inside terms in each set .. -3 * 3x = -9x last -- multiply last terms in each set .. -3 * 5 = -15 put them all together.. 12x^2 + 20x - 9x - 15 -- combine like terms 12x^2 + 11x - 15

OpenStudy (anonymous):

what about 4? and can u help me with a few more if u don't mind. ill go to your page and give u a medal on everything! please?

OpenStudy (texaschic101):

#4 is 6x^3 - 17h^2 - 11h + 2

OpenStudy (ria23):

haha, I only knew the third one.

OpenStudy (texaschic101):

I am not real good with ones like #4....I had to look that one up

OpenStudy (anonymous):

what is the factored form of the expression y^2+9y+8 a.(y+8)(y+1) B.(y-8)(y-1) C. (y+8)(y-1) D.(y-8)(y+1)**

OpenStudy (texaschic101):

a. (y + 8)(y + 1)

OpenStudy (texaschic101):

if you do FOIL on answer #a....then you will get y^2 + 9y + 8

OpenStudy (anonymous):

what is the factored form of the following expression x^2-4xy-21y^2 A.(x+3y)(x-7y)** B.(x-3y)(x-7y) C.(x+3y)(x+7y) D)x-3y)(x+7y)

OpenStudy (texaschic101):

you got this one right :)

OpenStudy (anonymous):

thank you soooo much!!! I would have failed without u!!!

OpenStudy (texaschic101):

the more problems you do like this, the better you will catch on. Be patient...keep practicing, and you will be the one helping people with this.

OpenStudy (anonymous):

can you help me with one more? this one im pretty sure but idk what is the degree of the monomial -2

OpenStudy (anonymous):

sorry!

OpenStudy (texaschic101):

the degree of the monomial is the highest exponent, if there is only 1, or it is the total of all the exponents, if there is more then one.

OpenStudy (anonymous):

its just -2

OpenStudy (texaschic101):

A monomial that is just a constant (a number with no variable) has a degree of 0

OpenStudy (texaschic101):

now if it is just a variable (a letter) then the degree is 1